The length and width of a rectangular swimming pool are each doubled, while the depth remains the same. Which is true?
a) it takes twice as much paint to paint the bottom
b) it takes twice as much paint to paint the four walls
c) it takes twice as much water to fill the pool
d) all of the above
e) none of the above

Answer by chriswen(106)

Lets say the original pool is l*w*h.
New pool is 2l*2w*h
a) the SA of bottom is the length times width.
Original Pool = l*w
New pool = 2l*2w = 4*l*w
FALSE
b) 4 walls is length times base times two. Added with width times base times two.
ORiginal Pool = l*b*2 + w*b*2
New Pool = 2l*b*2 + 2w*b*2 = 2(l*b*2 + w*b*2)
TRUE
c) Area of pool is length times width times base.
Original Pool = l*w*b
New Pool = 2l*2w*b = 4*l*w*b
FALSE
The correct answer is b)